OpenStudy (anonymous):

An unknown number y is 12 more than an unknown number x. The number y is also x less than 17. The equations to find x and y are shown below. y = x + 12 y = −x + 17 Which of the following statements is a correct step to find x and y? Multiply the equations to eliminate y. Write the points where the graphs of the equations intersect the x-axis. Write the points where the graphs of the equations intersect the y-axis. Add the equations to eliminate x.

OpenStudy (anonymous):

@aloud

OpenStudy (anonymous):

@badmood

OpenStudy (badmood):

Add the equations to eliminate x.

OpenStudy (anonymous):

u sure ?

OpenStudy (badmood):

y = 29/2

OpenStudy (anonymous):

true true

OpenStudy (badmood):

If you add those equations, you get: y+y=x+12−x+17

OpenStudy (anonymous):

can u help me wit like 3 more ?

OpenStudy (badmood):

sure

OpenStudy (anonymous):

Shiloh is 7 years older than Courtney. Shiloh's age is 13 years less than two times Courtney's age. The system below models the relationship between Shiloh's age (s) and Courtney's age (c): s = c + 7 s = 2c − 13 Which of the following methods is correct to find Shiloh's and Courtney's age? Solve s + 7 = 2s − 13 to find the value of s. Solve c + 7 = 2c − 13 to find the value of c. Write the points where the graphs of the equations intersect the x-axis. Write the points where the graphs of the equations intersect the y-axis.

OpenStudy (badmood):

I would say s = 2c − 13

OpenStudy (anonymous):

but which of the following methods is correct to find Shiloh's and Courtney's age?

OpenStudy (badmood):

Solve c + 7 = 2c − 13 to find the value of c.

OpenStudy (anonymous):

Variable a is 4 more than variable b. Variable a is also 1 less than b. Which of the following pairs of equations best models the relationship between a and b? a = b + 4 a = b − 1 a = 4b a = b + 1 b = a + 4 b = a − 1 b = 4a b = a + 1

OpenStudy (anonymous):

there in sets of 2 so theres four answers

OpenStudy (anonymous):

u still there ?